# GPU-Accelerated Algorithms for Process Mapping thanks: Henning Woydt is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – DFG SCHU 2567/6-1

### v1

@@ III-B Process Mapping
Similar to graph partitioning, process mapping has also seen a tremendous amount of research, and we refer the reader to Hoefler et. al [25] and Pellegrini [37] for broader overviews. Among the first to implement process mapping, was Hatazaki [22] who used graph partitioning to map MPI processes onto hardware topologies. Träff [46] and Yu et. al  [49] implemented mapping algorithms for special hardware, namely for the NEC SX-series of parallel vector computers and the Blue Gene/L Supercomputer, respectively.
@@
From a theoretical perspective it seems straightforward that minimizing $J(C,D,\Pi)$ will lead to a more performant algorithm. However, the questions remains if this is also true in practice. Brandfass et. al [7] showed that optimizing the communication cost indeed leads to performance improvements in the context of Computational Fluid Dynamics. Manwade et. al [33] show in a more general context that process mapping algorithms improve the performance of MPI applications.
@@
There are two main methods to solve process mapping, the two-phase approach and integrated mapping. The two-phase approach first leverages standard graph partitioning to determine a balanced $k$-way partition with small edge-cut of the task graph $G_{\mathcal{C}}$. The second phase then determines a 1-to-1 mapping of blocks to the PEs that minimizes the communication cost. Müller-Merbach [35] presented a greedy algorithm to obtain an initial mapping, that served as a basis for many future algorithms. After the mapping phase a potential third refinement phase can be used to determine individual vertex moves that decrease the communication cost. Heider [23] introduced a simple swap algorithm that considers all possible $\mathcal{O}(n^{2})$ swaps. This method was iteratively improved by Brandfass et. al [7] by avoiding redundant swaps, and Schulz and Träff [43], by utilizing better data structures and restricting the search space. Notable algorithms using the two-phase approach are GlobalMultisection [47] and SharedMap [44].
@@
GlobalMultisection [47] employs a method called hierarchical multisection inspired by KaFFPa-Map [43]. KaFFPa-Map would first create a balanced $k$-way partition and then create a communication model graph $G_{\mathbb{M}}$. This graph has $k$ vertices and edges connecting them if the corresponding $k$-way partition has edges between the blocks. $G_{\mathcal{M}}$ is then perfectly partitioned alongside the hierarchy $H=a_{1}:\ldots:a_{\ell}$, first into an $a_{\ell}$-way partition, then each of the blocks into an $a_{\ell-1}$-way partition and so forth. This is recursively repeated until $k$ blocks are obtained which represent the $k$ PEs of the system. The mapping of blocks to PEs, and therefore vertices to PEs, can be determined by following the recursive calls. GlobalMultisection improves the approach by not creating the communication model graph $G_{\mathbb{M}}$, but instead directly determining an $a_{\ell}$-way partition of $G_{\mathcal{C}}$ and then recursing on each block. Since no perfect balanced partition is needed, GlobalMultisection has more freedom to move vertices, resulting in lower communication costs.
@@
SharedMap [44] further decreases the communication cost. It introduces an adaptive imbalance strategy that rescales the imbalance parameter $\epsilon$ for each individual partitioning. It guarantees that the final $k$-way partition is $\epsilon$ balanced, and often gives each individual partitioning more space to work with. This results in an overall better mapping with lower communication costs. SharedMap uses KaFFPa [40] as one of its underlying graph partitioners, and it offers state-of-the-art solution qualities for process mapping.
@@
The second approach to solve the GPMP, integrated mapping, combines the partitioning and mapping into one algorithm. It leverages the multilevel scheme and the refinement algorithms minimize $J(C,D,\Pi)$. Two prominent algorithm are IntegratedMapping [15] and Mt-KaHyPar-Steiner [24].
@@
IntegratedMapping [15] uses matching based coarsening to create a smaller representation of the graph. The Global Path Algorithm [34], a $1/2$-approximation algorithm, is used to determine a matching of vertices. Instead of using the edge weights as a rating, a special rating function $\exp^{*}(\{u,v\})=\omega(\{u,v\})/c(u)c(v)$ from [26] is used. As initial partitioning, the hierarchical multisection approach is used. Refinement strategies include quotient graph refinement [40], $k$-way FM refinement [40], label propagation [38] and multi-try FM refinement [40]. All refinement methods use the gain of a vertex $v$ when moving it from its current block $\Pi(v)$ to a new block $b$. The gain is defined as follows:
$G_{b}(v)=\sum_{u\in N(v)}C_{v,u}\Big(D_{\Pi(v),\Pi(u)}-D_{b,\Pi(u)}\Big).$ (1)
@@
Multiple methods are presented to query the distance $D_{i,j}$ between two PEs $i$ and $j$. The simplest one saves a matrix with $\mathcal{O}(k^{2})$ space and access each distance in $O(1)$. The other methods, have a considerable lower space complexity, but come with more computational load.
@@
Mt-KaHyPar-Steiner [24] originally was developed to minimize wire lengths in VLSI designs. Logical units, hypergraph $\mathcal{H}$, are mapped to chip locations, edge-weighted graph $\mathcal{C}$, such that the total sum of used edge-weights, the implicit wire-length distance, is minimized. Replacing $\mathcal{H}$ with the task graph $G_{\mathcal{C}}$ and $\mathcal{C}$ with the hardware topology matrix $D$ leads to GPMP. Mt-KaHyPar-Steiner is integrated into the shared-memory parallel hypergraph partitioner Mt-KaHyPar [20]. The algorithm follows the multilevel (hyper-)graph partitioning approach. First the hypergraph is clustered, and these clustered are coarsed into a smaller graph. An initial $k$-way partition of the hypergraph is determined and then greedily mapped onto the chip locations. As local search refinement algorithms, label propagation [38], FM refinement [16] and flow-based refinement [40] are used.
III-C Kokkos
@@
We utilize the Kokkos [14] library to implement our algorithms. Kokkos is a C++ library providing performance portability across many-core architectures, including GPUs. It abstracts two main concepts: (1) Parallel execution patterns and (2) Multidimensional arrays. Since multidimensional arrays are not used in this work, we focus on the parallel execution patterns, which are built around three fundamental primitives. (a) parallel_for executes $N$ independent work units in parallel. Each work unit operates on the same user defined function $f(i)$ with only the index $i\in[0,N)$ changing for each work unit. (b) parallel_reduce performs a reduction $R=\bigoplus_{i=0}^{N-1}f(i)$ over values returned by $f(i)$ using an associative operator $\oplus$. (c) parallel_scan computes an exclusive prefix sum over a range of values i.e. $y_{i}=\sum_{j=0}^{i-1}f(j)$. All work units share access to memory, and Kokkos provides atomic operations and deterministic reduction mechanisms to prevent race conditions and ensure correctness on many-core architectures.
III-D CSR Graph Format
@@
A graph $G=(V,E)$ is typically stored in the Compressed Sparse Row (CSR) format (also referred to as Compressed Row Storage - CRS). The representation consists of an offset array $O$ (also called row-pointer) of size $|V|+1$ and two arrays of size $|E|$, one for the edge vertices $E_{v}$ and one for the edge weights $E_{w}$. The edges incident to vertex $v$ are stored in the contiguous range $[O[v],O[v+1]])$ of $E_{v}$ and $E_{w}$. Note that the graph is undirected and an edge $\{u,v\}$ is both stored for $u$ and $v$. This layout naturally supports parallelization across vertices, with each vertex processing its adjacency list independently.
IV GPU-Accelerated Process Mapping
The main contribution of this work is the development of two GPU-based approaches for solving the hierarchical process mapping problem. Given the close relationship between process mapping and graph partitioning, and the recent success of GPU-based partitioners in achieving faster runtimes, we aim to extend these advantages to process mapping.
@@ IV GPU-Accelerated Process Mapping
Our first algorithm follows the hierarchical multisection approach. It recursively partitions the task graph $G_{\mathcal{C}}$ along the systems hierarchy: first across islands, then racks, and so forth. The final mapping of blocks, and thus vertices, to PEs is obtained by following the recursive calls. The algorithm is presented in Section IV-A.
The second algorithm integrates process mapping directly into a multilevel scheme and is inspired by the GPU-based graph partitioner Jet [18] (Section III-A). Its refinement phase has been adapted to employ the new gain equation (Equation 1). Further details are given in Section IV-B.
IV-A Hierarchical Multisection
@@
The hierarchical multisection approach was first introduced by Kirchbach et al. [47]. It directly exploits the given hierarchy $H=a_{1}:\ldots:a_{\ell}$ of the supercomputer to partition the task graph $G_{\mathcal{C}}$. First, an $a_{\ell}$-way partition of $G_{\mathcal{C}}$ is determined utilizing a graph partitioning algorithm. The resulting $a_{\ell}$ blocks, represent the largest hardware components of the supercomputer (e.g., the islands). Since the graph partitioner minimizes the edge-cut between the blocks, it directly corresponds to minimizing the communication cost across the islands. Then the algorithm recursively partitions each of the $a_{\ell}$ blocks into an $a_{\ell-1}$-way partition. These blocks represent the second-largest component of the system (e.g., racks within islands). Again, minimizing the edge-cut corresponds to minimizing the communication across the racks of each island. This recursive procedure is continued until $k$ blocks are obtained. The mapping of the $k$ blocks to the $k$ PEs can be extracted by following the recursive calls.
@@
Schulz and Woydt [44] improved this method by introducing an adaptive imbalance parameter. Using the original imbalance parameter $\epsilon$ for each partitioning can overall result in an unbalanced $k$-way partition, i.e., some PEs may be overloaded. To avoid this problem, an adaptive imbalance parameter $\epsilon^{\prime}$ is determined for each partitioning. It depends on the original imbalance $\epsilon$, the overall graph weight $c(V)$, the weight of the current subgraph $c(V^{\prime})$, the desired number of blocks $k$, the number of total blocks the current subgraph will be partitioned into $k^{\prime}$ and the depth of the current partition in the hierarchy $d$. The adaptive imbalance is then determined by
@@
Using the adaptive imbalance guarantees that the final partition is $\epsilon$-balanced. Experimental results in [44] demonstrate that the adaptive imbalance improves the overall mapping quality. In particular, the SharedMap [44] implementation currently achieves state-of-the-art solution quality for the HPMP.
To adopt the hierarchical multisection approach to GPUs, we must (1) employ a GPU-based graph partitioner and (2) construct the resulting subgraphs directly on the GPU.
@@
For the partitioning step, we use Jet [18] which we already presented in Section III-A. Jet outperforms CPU-based graph partitioners in terms of speed by factor in the range from 2 to 20 times while delivering equal or better solution quality. Alternative GPU-based graph partitioners such as Sphynx [1] or G-kway [31] could also be used. As no comprehensive comparison between GPU partitioners exists, we adopt Jet for its thorough evaluation. However, any faster or higher-quality alternative would translate to our algorithm.
For the second step, we need to build the subgraphs of a graph $G=(V,E)$ given a partition $\Pi:[n]\mapsto[k]$. To circumvent the costly down- and upload between the CPU and GPU, the entire process is executed on the GPU. Our algorithm works in $k$ iterations, where each iteration builds one subgraph $G_{i}$. Each iteration consists of three phases. The first phase determines the number of vertices $n_{i}$, edges and the subgraphs weight using three parallel_reduce loops. A vertex $v$ belongs to the $i$th subgraph if $\Pi(v)=i$ and likewise an edge $\{u,v\}$ belongs to it, if $\Pi(u)=\Pi(v)=i$. The second phase determines a mapping between the vertices of $G_{i}$ to the vertices of $G_{\mathcal{C}}$. The vertices of $G_{i}$ need to be indexed from 0 to $n_{i}-1$, which is also incorporated into the mappings. The mappings are needed so we can propagate the assigned partition of vertices in $G_{i}$ back to the vertices of $G_{\mathcal{C}}$. This step is efficiently implemented via a parallel_scan over the vertex set. The prefix sum assigns each vertex a unique new vertex in the range [0,n1)[0,n-1). In the third phase, we construct the subgraph using a parallel_scan. The scan first computes the CSR offsets for each vertex and then writes the corresponding edges and weights. For each vertex $u$ only edges $\{u,v\}$ in the same subgraph are inserted, while the mapping functions translate between global and local vertex IDs. Subgraph creation usually takes less than 2% of the total runtime of the complete algorithm.
Algorithm 1 presents the pseudocode of the hierarchical multisection method. The function takes as input a (sub)graph $G_{L}$, the hierarchy $H$, the current hierarchy level $i$, an identifier list $I$, and a local mapping $\Pi_{L}$ that maps the vertices of $G_{L}$ to the vertices of $G_{\mathcal{C}}$. The identifier list $I$ uniquely identifies the position of the subgraph within the system hierarchy. Lines 242-4 handle the exit condition: when $i=0$, the subgraph represents one of the final $k$ blocks and all its vertices are assigned to exactly one of the $k$ PEs. The function $\textsc{calcID}(H,I)$ determines a unique ID for each identifier $I$ such that neighboring components receive consecutive IDs. The mapping $\Pi_{L}$ is used in Line 5 to translate the local vertices back to their original vertices in $G_{\mathcal{C}}$ and for these vertices the mapping is set to the ID.
If the subgraph $G_{L}$ resides on a higher level ($i>0$) then further partitioning is required (Lines 6126-12). First, The adaptive imbalance $\epsilon^{\prime}$ is determined and then Jet [18] is utilized to partition $G_{L}$ into $a_{i}$ subgraphs via the mapping $\Pi^{\prime}:V(G_{L})\mapsto[a_{i}]$. For each resulting subgraph $G^{\prime}_{L}$, we determine a local mapping $\Pi^{\prime}_{L}$ from its vertices to $G_{\mathcal{C}}$ (needed in Line 5). The identifier list $I$ is copied and extended by index $j$ to uniquely identifier the subgraph (needed in Line 3). The algorithm then recurses on each subgraph. Finally, the complete algorithm is invoked in Line 14 with the task graph $G_{\mathcal{C}}$, the hierarchy $H$, the highest hierarchy level $\ell$, an empty identifier list, and the identity mapping for $G_{\mathcal{C}}$.
@@
1:Graph $G$, Hierarchy $\smash{H=a_{1}:\ldots:a_{\ell}}$
@@
3:  if $i=0$ then
4:   $\textsc{ID}\leftarrow\textsc{calcID}(H,I)$
5:   for $\smash{v\in V_{L}}$ do $\smash{\Pi(\Pi_{L}(v))\leftarrow\textsc{ID}}$
6:  $\smash{\epsilon^{\prime}\leftarrow\textsc{AdapImb}(\ldots)}$ $\triangleright$ uses Eq. 2
7:  $\smash{\Pi^{\prime}\leftarrow\textsc{Jet}(G_{L},a_{i},\epsilon^{\prime})}$
8:  $j\leftarrow 0$
9:  for $G^{\prime}_{L}\in\textsc{CreateSubgraphs}(G_{L},\Pi^{\prime})$ do
10:   $\Pi^{\prime}_{L}\leftarrow\textsc{MapToG}(G^{\prime}_{L},\Pi_{L})$
11:   $\smash{\textsc{HM}(G^{\prime}_{L},H,i-1,I+[j],\Pi^{\prime}_{L})}$
12:   $\smash{j\leftarrow j+1}$
13:end function
[…]
IV-B Integrated Mapping

### v2

@@ 3.2 Process Mapping
Similar to graph partitioning, process mapping has also seen a tremendous amount of research, and we refer the reader to Hoefler et al. [26] and Pellegrini [38] for broader overviews. Among the first to implement process mapping was Hatazaki [23] who used graph partitioning to map MPI processes onto hardware topologies. Träff [47] and Yu et al. [50] implemented mapping algorithms for special hardware, namely for the NEC SX-series of parallel vector computers and the Blue Gene/L Supercomputer, respectively.
@@
From a theoretical perspective, it appears straightforward that minimizing $J(\mathcal{C},\mathcal{D},\Pi)$ will lead to a more efficient algorithm. However, the question remains whether the theoretical improvement translates into practical performance gains. Brandfass et al. [8] show that optimizing the communication cost indeed leads to performance improvements for the simulation of Computational Fluid Dynamics. Manwade et al. [34] show in a more general context that process mapping algorithms improve the performance of MPI applications.
@@
There are two main methods to solve process mapping: the two-phase approach and integrated mapping. The two-phase approach first leverages standard graph partitioning to determine a balanced $k$-way partition with small edge-cut of the task graph $G_{\mathcal{C}}$. This results in a partition with an overall small communication volume between the blocks. The second phase then determines a 1-to-1 mapping of blocks to the PEs that minimizes the communication cost. Müller-Merbach [36] presented a greedy algorithm to obtain an initial mapping that served as a basis for many future algorithms. After the mapping phase, a potential third refinement phase can be used to determine individual vertex moves that decrease the communication cost. Heider [24] introduced a simple swap algorithm that considers all possible $\mathcal{O}(n^{2})$ swaps. This method was iteratively improved by Brandfass et al. [8] by avoiding redundant swaps, and Schulz and Träff [44] by utilizing better data structures and restricting the search space. Notable algorithms are GlobalMultisection [48] and SharedMap [45].
@@
GlobalMultisection [48] employs a method called hierarchical multisection inspired by KaFFPa-Map [44]. KaFFPa-Map would first create a balanced $k$-way partition and then create a communication model graph $G_{\mathbb{M}}$. This graph has $k$ vertices and edges connecting them if the corresponding $k$-way partition has edges between the blocks. $G_{\mathcal{M}}$ is then perfectly partitioned alongside the hierarchy $H=a_{1}:\ldots:a_{\ell}$, first into an $a_{\ell}$-way partition, then each of the blocks into an $a_{\ell-1}$-way partition and so forth. This is repeated recursively until $k$ blocks are obtained. These represent the $k$ PEs of the system. The mapping of blocks to PEs, and therefore vertices to PEs, can be determined by following the recursive calls. GlobalMultisection improves the approach by not creating the communication model graph $G_{\mathbb{M}}$, but instead directly determining an $a_{\ell}$-way partition of $G_{\mathcal{C}}$ and then recursing on each block. Since no perfectly balanced partition is needed, GlobalMultisection has more freedom to move vertices, resulting in lower communication costs.
@@
SharedMap [45] further decreases the communication cost. It introduces an adaptive imbalance strategy that rescales the imbalance parameter $\epsilon$ for each individual partitioning. It guarantees that the final $k$-way partition is $\epsilon$-balanced. This results in an overall better mapping with lower communication costs. SharedMap uses the serial graph partitioner KaFFPa [41] as one of its underlying graph partitioners, and currently offers state-of-the-art solution quality for HPMP.
@@
The second approach to solve the GPMP, integrated mapping, combines the partitioning and mapping into one algorithm. It leverages the multilevel graph partitioning scheme and the refinement algorithms minimize $J(\mathcal{C},\mathcal{D},\Pi)$ instead of edge-cut. Two prominent algorithms are IntMap [16] and Mt-KaHyPar-Steiner [25].
@@
IntMap [16] uses matching-based coarsening to create a smaller representation of the graph. The Global Path Algorithm [35], a $1/2$-approximation algorithm, is used to determine a matching of vertices. Instead of using the edge weights as a rating, a special rating function $\exp^{*}(\{u,v\})=\omega(\{u,v\})/c(u)c(v)$ from [27] is used. As initial partitioning, the hierarchical multisection approach is used. Refinement strategies include quotient graph refinement [41], $k$-way FM refinement [41], label propagation [39], and multi-try FM refinement [41]. All refinement methods use the gain of a vertex $v$ when moving it from its current block $\Pi(v)$ to a new block $b$. The gain is defined as follows:
$G_{b}(v)=\sum_{u\in N(v)}C_{v,u}\Big(\mathcal{D}_{\Pi(v),\Pi(u)}-\mathcal{D}_{b,\Pi(u)}\Big).$ (1)
@@
Multiple methods are presented to query the distance $\mathcal{D}_{i,j}$ between two PEs. The simplest one saves a matrix with $\mathcal{O}(k^{2})$ space and accesses each distance in $O(1)$. The other methods, have a considerably lower space complexity, but come with more computational load.
@@
Mt-KaHyPar-Steiner [25] originally minimizes wire lengths in VLSI designs. Logical units, a hypergraph $\mathcal{H}$, are mapped to chip locations, an edge-weighted graph $\mathcal{C}$, such that the total sum of used edge weights, the implicit wire-length distance, is minimized. Replacing $\mathcal{H}$ with the task graph $G_{\mathcal{C}}$ and $\mathcal{C}$ with the hardware topology matrix $\mathcal{D}$ leads to GPMP. Mt-KaHyPar-Steiner is integrated into the shared-memory parallel hypergraph partitioner Mt-KaHyPar [21]. The algorithm follows the multilevel (hyper-)graph partitioning approach. First, the hypergraph is clustered, and these clusters are contracted into individual vertices, creating a smaller graph. An initial $k$-way partition of the hypergraph is determined and then greedily mapped onto the chip locations. As local search refinement algorithms, label propagation [39], FM refinement [17], and flow-based refinement [41] are used.
3.3 Kokkos
@@
We utilize Kokkos [15], a C++ library providing performance portability across many-core architectures, including GPUs. It abstracts two main concepts: (1) parallel execution patterns and (2) multidimensional arrays. Since multidimensional arrays are not used in this work, we focus on the parallel execution patterns, which are built around three fundamental primitives. (a) parallel_for executes $N$ independent work units in parallel. Each work unit operates on the same user-defined function $f(i)$ with only the index $i\in[0,N)$ changing for each work unit. (b) parallel_reduce performs a reduction $R=\bigoplus_{i=0}^{N-1}f(i)$ over values returned by $f(i)$ using an associative operator $\oplus$. (c) parallel_scan computes an exclusive prefix sum over a range of values, i.e. $y_{i}=\sum_{j=0}^{i-1}f(j)$.
3.4 CSR Graph Format
@@
A graph $G=(V,E)$ is typically stored in the Compressed Sparse Row (CSR) format. The representation consists of an offset array $O$ (also called row-pointer) of size $|V|+1$ and two arrays of size $|E|$, one for the edge vertices $E_{v}$ and one for the edge weights $E_{w}$. The edges incident to vertex $v$ and the corresponding edge weights are stored in the contiguous range $[O[v],O[v+1])$ of $E_{v}$ and $E_{w}$. Note that the graph is undirected and an edge $\{u,v\}$ is stored for both $u$ and $v$. This layout naturally supports parallelization across vertices.
4 GPU-Accelerated Process Mapping
@@
The main contribution of this work is the development of two GPU-based approaches for solving the hierarchical process mapping problem. Given the close relationship between process mapping and graph partitioning, and the recent success of GPU-based partitioners in achieving faster runtimes, we extend these advantages to process mapping. The first algorithm follows the hierarchical multisection approach. It recursively partitions the task graph $G_{\mathcal{C}}$ along the system’s hierarchy: first across the islands, then across the racks, and so forth. The final mapping of blocks, and thus vertices, to PEs is obtained by following the recursive calls. The algorithm is presented in Section 4.1. The second algorithm integrates process mapping directly into the multilevel graph partitioning scheme and is inspired by the GPU-based graph partitioner Jet [19]. Its refinement phase has been adapted to employ the new gain equation (Equation 1). Section 4.2 presents the algorithm.
Extended CSR Format. Before presenting the algorithms, we want to present our extended CSR format. As introduced in Section 3.4, the CSR representation naturally supports parallelization over vertices, allowing each thread to process a vertex’s adjacency list independently. However, some computations benefit from direct parallelization over edges. In standard CSR, this requires nested parallel loops, first over vertices and then their neighbors. This can lead to poor load balancing due to varying vertex degrees, resulting in idle threads on GPUs. To overcome this, we extend the CSR format with an additional array $E_{u}$ of size $|E|$ that explicitly stores the opposite endpoint of each edge. This edge-centric representation enables flat parallelism over all edges, improving load balance and GPU utilization. We refer to this edge-list representation as $\mathbb{E}$ throughout the rest of this work.
4.1 Hierarchical Multisection
@@
The hierarchical multisection approach was first introduced by Kirchbach et al. [48]. It directly exploits the given hierarchy $H=a_{1}:\ldots:a_{\ell}$ of the supercomputer to partition the task graph $G_{\mathcal{C}}$. First, an $a_{\ell}$-way partition of $G_{\mathcal{C}}$ is determined utilizing a graph partitioning algorithm. The resulting $a_{\ell}$ blocks represent the largest hardware components of the supercomputer (e.g., the islands). Since the graph partitioner minimizes the edge-cut between the blocks, it directly corresponds to minimizing the communication cost across the islands. Then the algorithm recursively partitions each of the $a_{\ell}$ blocks into an $a_{\ell-1}$-way partition. These blocks represent the second-largest component of the system (e.g., racks within islands). Again, minimizing the edge-cut corresponds to minimizing the communication across the racks of each island. This recursive procedure is continued until $k$ blocks are obtained.
@@
Schulz and Woydt [45] improved this method by introducing an adaptive imbalance parameter. Using the original imbalance parameter $\epsilon$ for each partitioning can result overall in an unbalanced $k$-way partition. To avoid this problem, an adaptive imbalance parameter $\epsilon^{\prime}$ is determined for each partitioning. It depends on the original imbalance $\epsilon$, the overall graph weight $c(V)$, the weight of the current subgraph $c(V^{\prime})$, the desired number of blocks $k$, the number of total blocks the current subgraph will be partitioned into $k^{\prime}$ and the depth of the current partition in the hierarchy $d$. The adaptive imbalance is then determined by
@@
Using the adaptive imbalance guarantees that the final partition is $\epsilon$-balanced. Experimental results in [45] demonstrate that the adaptive imbalance improves the overall mapping quality. In particular, SharedMap [45] currently achieves state-of-the-art quality for the HPMP.
@@
To adopt the hierarchical multisection approach to GPUs, we must (1) employ a GPU-based graph partitioner to obtain the $a_{i}$-way partitions and (2) construct the resulting subgraphs directly on the GPU. For the partitioning step, we use our own reimplementation of Jet [19], which we already presented in Section 3.1. Jet outperforms the CPU-based graph partitioner Metis [28] in terms of speed by a factor in the range of 2 to 20 times while delivering equal or better solution quality. Alternative GPU-based graph partitioners such as Sphynx [1] or G-kway [32] could also be used. However, no comprehensive comparison between GPU partitioners exists, so we adopt the implementation of Jet for its thorough evaluation. Any faster or higher-quality graph partitioner would also improve this algorithm.
For the second step, the induced subgraphs of a graph $G=(V,E)$ given a partition $\Pi:[n]\mapsto[k]$ need to be built. To circumvent the costly download and upload between the CPU and GPU, the entire process is executed on the GPU. The algorithm works in $k$ iterations, where each iteration builds one subgraph $G^{\prime}$. Algorithm 1 shows the pseudocode of one loop iteration. We define
$\mathbf{1}[P]=\begin{cases}1&\text{if }P\text{ is true}\\ 0&\text{otherwise}\end{cases}$
as an indicator function, which returns $1$ if the predicate $P$ is true and 0 else.
Each iteration consists of three phases. The first phase determines the number of vertices $n^{\prime}$, edges $m^{\prime}$ and the subgraph’s weight $w^{\prime}$ using three parallel_reduce loops. A vertex $v$ belongs to the subgraph with id $k^{\prime}$ if $\Pi(v)=k^{\prime}$ and likewise an edge $\{u,v\}$ belongs to it if $\Pi(u)=\Pi(v)=k^{\prime}$. Lines 1 – 3 show this phase. The second phase remaps the vertices $v\in V_{i}$ from $[n]$ to $n^{\prime}]$. This remapping is a technical requirement, since graph partitioners expect a continuous range of vertices. The mapping $M:[n]\mapsto[n^{\prime}]$ is also later needed to propagate the resulting partition to the original vertices. Lines 585-8 show the pseudocode. Note that in Kokkos, a single parallel_scan is sufficient to construct the mapping. In the third phase, the subgraph is constructed. Lines 10 to 26 show the process. At first the new degree of each vertex is determined in parallel (Lines 10–13). Based on the new degrees the row pointer array of the graph can be determined using a prefix sum (Line 12). This assigns each vertex in the subgraph a continuous range, to place its edges. The variables $E^{\prime}$ in Lines 15–17 are the corresponding arrays of the subgraph. At last the actual insertion of the edges into the subgraph is done in Lines 18–26. For each vertex $v$ that belongs to the subgraph, its neighborhood is iterated over. For each neighbor $u$ that also belongs to the subgraph, the edge is inserted into the arrays and the index $i$ is incremented. Note that this whole phase can also be implemented as one large parallel_scan. Subgraph creation usually takes less than 5% of the total runtime of the hierarchical multisection algorithm.
Algorithm 1 Build Subgraph
1:Graph $G$, Partition $\Pi$, Subgraph ID $k^{\prime}$
[…]
5:
[…]
7:for $v\in V$ do in parallel
8:  $B[v]\leftarrow\textbf{1}[\Pi(v)=k^{\prime}]$
[…]
10:
[…]
12:for $v\in V$ do in parallel
13:  if $\Pi(v)=k^{\prime}$ then
14:   $D[M(v)]\leftarrow\sum_{u\in N(v)}\textbf{1}[\Pi(u)=k^{\prime}]$
[…]
19:for $v\in V$ do in parallel
20:  if $\Pi(v)=k^{\prime}$ then
21:   $i\leftarrow R^{\prime}[M(v)]$
22:   for $u\in N(v)$ do
23:     if $\Pi(u)=k^{\prime}$ then
24:      $E^{\prime}_{v}[i]\leftarrow M(u)$
25:      $E^{\prime}_{w}[i]\leftarrow\omega(u,v)$
26:      $E^{\prime}_{u}[i]\leftarrow M(v)$
27:      $i\leftarrow i+1$
28:return $G^{\prime}=(R^{\prime},E^{\prime}_{v},E^{\prime}_{w},E^{\prime}_{u}),\ M$
The complete hierarchical multisection algorithm can be seen in Algorithm 2. Here, it is implemented as a recursive function but an iterative approach would also be possible. The function takes as input a (sub)graph $G^{\prime}$, the hierarchy $H$ and the current hierarchy level $i$. Lines 2 and 3 handle the partitioning of the graph. At first the adaptive imbalance $\epsilon^{\prime}$ is computed using Equation 2 and afterwards our reimplementation of Jet [19] is used to partition the graph on the GPU. It determines a partition $\Pi^{\prime}$ into $a_{i}$ blocks with the specified $\epsilon^{\prime}$ imbalance. If this was the last partitioning of this recursive call ($i=1$) then the determined partition $\Pi^{\prime}$ needs to be propagated to the actual partition $\Pi$ of $G$. Since this is only a technical process we omit its details and only outline the brief idea. For each created graph $G^{\prime}$ keep a mapping $M^{\prime}$ from $V^{\prime}$ to the original vertices $V$ in $G$. Then for each $v\in V^{\prime}$ set $\Pi(M^{\prime}(v))=\Pi^{\prime}(v)$ into the final partition. The mapping $M^{\prime}$ can be computed by utilizing the mapping $M$ from Algorithm 1. If the subgraph $G^{\prime}$ resides on a higher level ($i>1$), then the resulting subgraphs need to be constructed and further partitioned. In Line 7 the subgraphs are built using Algorithm 1. Each subgraph is then recursively passed to the function, but with i1i-1 it is indicated that the subgraphs reside one level further down in the hierarchy. Finally, the complete algorithm is invoked in Line 10 with the task graph $G_{\mathcal{C}}$, the hierarchy $H$ and the highest hierarchy level $\ell$.
@@
1:Graph $G$, Hierarchy $H=a_{1}:\ldots:a_{\ell}$, Imbalance $\epsilon$
2:function HM($G^{\prime},H,i$)
3:  $\epsilon^{\prime}\leftarrow\textsc{AdapImb}(\ldots)$ $\triangleright$ uses Equation 2
4:  $\Pi^{\prime}\leftarrow\textsc{GPU-GP}(G^{\prime},a_{i},\epsilon^{\prime})$ $\triangleright$ GPU Graph Partitioner
5:  if $i=1$ then
6:   propagate $\Pi^{\prime}$ to final partition $\Pi$
7:   return
8:  for $G^{\prime\prime}\in\textsc{CreateSubgraphs}(G^{\prime},\Pi^{\prime})$ do $\triangleright$ uses Algorithm 1
9:   $\textsc{HM}(G^{\prime\prime},H,i-1)$}
10:end function
@@
4.2 Integrated Mapping